OpenStudy (anonymous):

2. Using Shell Method find the volume of the solid of revolution generated by revolving the following area: (a) y = x , y = -x/2 , x = 2 : about y-axis

OpenStudy (sirm3d):

\[\mathrm dV=2\pi x \left(x-\frac{-x}{2}\right) \mathrm dx\]